OpenStudy (anonymous):

Simplify by removing factors of 1 q^2-49 over (q+7)^2

OpenStudy (anonymous):

\[\left(\begin{matrix}q^2-49 \\ (q+7)^2\end{matrix}\right)\]

OpenStudy (anonymous):

Btw, if you want do fractions bad, you can type frac{top}{bottom} to get: \[\frac{top}{bottom}\] Anyways, factor the top giving: \[\frac{(q+7)(q-7)}{(q+7)^2}=\frac{q-7}{q+7}\]

OpenStudy (anonymous):

ok so it's q-7/q+7? Sorry I don't understand algebra

OpenStudy (anonymous):

Well you can "cancel" the common factor. In this case, q+7. So yes.

OpenStudy (anonymous):

For example, if you had: \[\frac{4}{2}=\frac{2*2}{2}=\frac{2}{1}=2\] Its the same concept.

OpenStudy (anonymous):

Oh ok I see it now! Thank you

OpenStudy (anonymous):

No problem :P